Students collected 600 cans for a canned food drive. That was 80% of their goal. How many more kids do they need to collect to reach their goal?
1
Expert's answer
2019-09-17T11:23:26-0400

600=0.8x

x=600/0.8=750

750-600=150.

Student needs to collect 150 more cans.